OpenStudy (anonymous):

What is the simplified form of x^2 - 36/ x +6 i got x-6 with the restriction x= -6

OpenStudy (anonymous):

x^2-36 can be written as (x-6)*(x+6). Since a^2 - b^2 = (a+b)*(a-b). On simplifying, the answer is x-6.

OpenStudy (anonymous):

And in this case a= x and b = 6.

OpenStudy (anonymous):

so i was right ? :-) i know it ='s x-6 but would the restriction be -6 or 6 thats the part im confused about

OpenStudy (anonymous):

The restriction for (x^2 - 36)/ x +6 is x=-6. Since if x was -6 the denominator will become zero and the value of the expression will explode.

OpenStudy (anonymous):

when you factor x^2-36 u find= (x+6)(x-6) and then u'll have (x+6)(x-6)/x+6 => x-6

OpenStudy (anonymous):

yess i was right :-)

OpenStudy (anonymous):

Yes, congratulations :) .

OpenStudy (anonymous):

applause*